OpenStudy (moneybaby980):

The price of a video game is $32.99 before tax. Andre bought the video game for 20% off. He then paid 8% sales tax on the discounted price. Part A: How much did Andre pay in sales tax? Show all work and steps in your solution. Part B: What is the total amount that Andre paid for the video game? Show all work and steps in your solution.

OpenStudy (anonymous):

\[32.99*0.2\] will get you the 20% off

OpenStudy (anonymous):

subtract that from 32.99

OpenStudy (anonymous):

\[32.99-6.60=26.39\]

OpenStudy (moneybaby980):

ok

OpenStudy (anonymous):

that's your amount without sales tax

OpenStudy (anonymous):

\[26.39*0.08=2.11\] 2.11 is your sales tax

OpenStudy (anonymous):

theres part A

OpenStudy (anonymous):

do all that for part B just add 26.39+2.11=28.5

OpenStudy (moneybaby980):

y do u add

OpenStudy (anonymous):

to get the full amount you paid